# Collaborative Learning with Different Labeling Functions

### v2

@@ Abstract
We show that, when the data distributions satisfy a weaker realizability assumption, sample-efficient learning is still feasible. We give a learning algorithm based on Empirical Risk Minimization (ERM) on a natural augmentation of the hypothesis class, and the analysis relies on an upper bound on the VC dimension of this augmented class.
@@ 1 Introduction
  • We formalize a model of collaborative learning with different labeling functions, and a sufficient condition, termed $(k,\epsilon)$-realizability, for sample-efficient collaborative learning. Under this assumption, we give a learning algorithm with sample complexity $O(kd\log(n/k)+n\log n)$. This algorithm is based on Empirical Risk Minimizarion (ERM) over an augmentation of the hypothesis class.
@@ 1.2 Our Results
Distributions $\mathcal{D}_{1}$, $\mathcal{D}_{2}$, normal-$\ldots$, $\mathcal{D}_{n}$ are $(k,\epsilon)$-realizable with respect to hypothesis class $\mathcal{F}$, if there exist $f^{*}_{1},f^{*}_{2},\ldots,f^{*}_{k}\in\mathcal{F}$ such that $\min_{j\in[k]}\mathrm{err}_{\mathcal{D}_{i}}(f^{*}_{j})\leq\epsilon$ holds for every $i\in[n]$.
@@
Suppose that $\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{n}$ are $(k,0)$-realizable with respect to a $2$-refutable hypothesis class $\mathcal{F}$. Then, for any $\delta>0$, there is an $(\epsilon,\delta)$-PAC algorithm that runs in $\mathrm{poly}(n,k,1/\epsilon,\log(1/\delta))$ time, makes $\mathrm{poly}(n)$ calls to an ERM oracle for $\mathcal{F}$, and has a sample complexity of
@@ 1.3 Related Work
Most closely related to our work are the previous studies of Collaborative PAC Learning [BHPQ17, NZ18, CZZ18] and related fields such as multi-task learning [HK22], multi-distribution learning [HJZ22, AHZ23, Pen23, ${}^{+}$23], federated learning [${}^{+}$17, MSS19, CCD23] and multi-source domain adaptation [MMR08, KL19, ${}^{+}$21].
In multi-task learning/multi-source domain adaptation, there are $n$ distributions, each with a fixed number of samples, and our goal is to use these samples to learn a hypothesis that has a small risk on some target distribution. A line of works [${}^{+}$10, KL19, ${}^{+}$21] studied the generalization risk in this scenario, but their bounds all depend on the discrepancy among $n$ distributions and contain non-vanishing residual constants. To avoid this residual constant in the bounds, [HK22] considered a Bernstein condition assumption on the hypothesis class and some transferrability assumptions between the $n$ source distributions and the target distribution. They studied the minimax rate of this learning scenario, and gave a nearly-optimal adaptation algorithm. Specially, some works studied the multi-task linear regression [${}^{+}$20, ${}^{+}$23]. [${}^{+}$20] considered linear regression from multiple distributions, where all tasks share the same input feature covariate $\mathbf{X}\in\mathbb{R}^{m\times d}$, but with different labeling functions. They designed the Hard Parameter Sharing estimator and established an excess risk upper bound of $O(\frac{d\sigma^{2}}{m}+\mathrm{heterogeneity})$. Recently, [${}^{+}$23] proposed the $s$-sparse heterogeneity assumption among the labeling functions in multi-task linear regression, and designed an algorithm which achieves an $O(\frac{s\sigma^{2}}{m_{i}}+\frac{d\sigma^{2}}{\sum_{i=1}^{n}m_{i}})$ excess risk on the $i$-th task. Notice that when $s$ is smaller than $d$, this bound is strictly sharper than individual learning bound $O(\frac{d\sigma^{2}}{m_{i}})$. The difference between our scenario and multi-task learning is that we allow the learner to draw an arbitrary number of samples from each distribution, instead of assuming that each distribution only has a fixed number of samples.
@@
Multi-distribution learning was recently proposed by [HJZ22], where $n$ players try to learn a single model $\hat{f}$, that can have an $\epsilon$ excess error on the worst case distribution among $n$ players, i.e., $\max_{i\in[n]}\mathrm{err}_{\mathcal{D}_{i}}(\hat{f})\leq\epsilon+\mathrm{OPT}$. They gave an algorithm with sample complexity $O(\frac{d\log n}{\epsilon^{2}}+\frac{nd\log(d/\epsilon)}{\epsilon})$, and proved a lower bound of $\tilde{\Omega}(\frac{d+k}{\epsilon^{2}})$. Note that this is a more pessimistic learning guarantee than ours, since the value $\mathrm{OPT}$ can be very large. Very recently, two concurrent papers [Pen23, ${}^{+}$23] gave algorithms that match this lower bound, resolving some of the open problems formulated in [AHZ23].
@@
Another recent line of work [${}^{+}$20, KSKO20, DJKS23, ${}^{+}$23] studied learning mixtures of linear regressions from data batches. In these setups, there are $k$ unknown linear regression models. Each data batch consists of labeled examples produced by one of the linear models chosen randomly. This line of work gave trade-offs between the number of batches and the batch size in order for the parameters of the $k$ linear models to be efficiently learnable.
In comparison, our model allows the learner to adaptively sample from the data distributions, whereas the batch sizes are fixed in the model of learning with batches. We also note that the results for learning with batches require assumptions on the marginal distribution, such as Gaussianity or certain hypercontractivity and condition number properties. Also, except the recent work of [${}^{+}$23], they all required the marginal distribution of the instance to be the same across all batches.
@@ 3 Sample Complexity Upper Bound
Lemma 11 implies that in general, $\mathcal{F}_{G,k}$ has a VC dimension of $O(kd+|G|\log k)$. To gain some intuition behind the bound in Lemma 11, suppose that $\mathcal{F}$ is a finite class of size $2^{d}$. By Definition 2, the size of $\mathcal{F}_{n,k}$ is at most $|\mathcal{F}|^{k}\cdot k^{n}=2^{kd+k\log_{2}n}$, and the bound in the lemma immediately follows. The actual proof, of course, needs to deal with the case that $\mathcal{F}$ is larger or even infinite.
@@ 5.2 Hardness under Two Labeling Functions
We first reduce the general subset sum problem to a special case, in which the $n$ numbers sum up to $2^{n+1}$ and the target value is exactly $2^{n}$. Let ($(\{a_{i}\}_{i\in[m]},t)$ be an instance of subset sum (i.e., deciding whether there exists $T\subseteq[m]$ such that $\sum_{i\in T}a_{i}=t$). Let $s=\sum_{i=1}^{m}a_{i}$ and pick $n=\max\{m+2,\lfloor\log_{2}s\rfloor+1\}$ such that $n-m\geq 2$ and $2^{n}>s$. We pad $n-m$ numbers to the instance, such that $a_{m+1}=a_{m+2}=\cdots=a_{n-2}=0$, $a_{n-1}=2^{n}-t$, $a_{n}=2^{n}-(s-t)$. Note that the size of the instance increases at most polynomially after the padding. Also note that the numbers now sum up to
@@ 6 An Efficient Algorithm for Identical Marginals
In this section, we prove Theorem 4, which addresses the special case that $\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{n}$ share the same marginal distribution over $\mathcal{X}$. In this case, we show that there is a simple algorithm that efficiently clusters the distributions and learn accurate classifiers using $\ll nd$ samples. The algorithm is formally defined as Algorithm 1.

### v3

@@ Abstract
We show that, when the data distributions satisfy a weaker realizability assumption, which appeared in [CM12] in the context of multi-task learning, sample-efficient learning is still feasible. We give a learning algorithm based on Empirical Risk Minimization (ERM) on a natural augmentation of the hypothesis class, and the analysis relies on an upper bound on the VC dimension of this augmented class.
@@ 1 Introduction
  • We formalize a model of collaborative learning with different labeling functions, and a sufficient condition, termed $(k,\epsilon)$-realizability, for sample-efficient collaborative learning. This realizability assumption was used by [CM12] in the context of multi-task learning. Under this assumption, we give a learning algorithm with sample complexity $O(kd\log(n/k)+n\log n)$. This algorithm is based on Empirical Risk Minimizarion (ERM) over an augmentation of the hypothesis class.
@@ 1.2 Our Results
Distributions $\mathcal{D}_{1}$, $\mathcal{D}_{2}$, $\ldots$, $\mathcal{D}_{n}$ are $(k,\epsilon)$-realizable with respect to hypothesis class $\mathcal{F}$, if there exist $f^{*}_{1},f^{*}_{2},\ldots,f^{*}_{k}\in\mathcal{F}$ such that $\min_{j\in[k]}\mathrm{err}_{\mathcal{D}_{i}}(f^{*}_{j})\leq\epsilon$ holds for every $i\in[n]$.
@@
While we prove Theorem 1 (as well as the other positive results in the paper) under the assumption that the learning algorithm is given the value of $k$, this assumption can be removed via a standard doubling trick: We consider a sequence of guesses on the value of $k$: $1=k_{1}<k_{2}<k_{3}<\cdots$, where each $k_{i+1}$ is the smallest value of $k$ such that the sample complexity bound is at least twice the bound for $k_{i}$. Then, we run the learning algorithm with $k$ set to $k_{1},k_{2}-1,k_{2},k_{3}-1,k_{3},\ldots$ in order, and terminate the algorithm as soon as we are convinced that the actual $k$ is larger than the current guess. This procedure succeeds as soon as the guess exceeds the actual value of $k$, and the sample complexity only increases by a constant factor.
@@
Our algorithm for the theorem above follows a similar approach to the lifelong learning algorithms of [BBV15, PU16].
@@
Suppose that $\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{n}$ are $(k,0)$-realizable with respect to a $2$-refutable hypothesis class $\mathcal{F}$. For any $\delta>0$, there is an $(\epsilon,\delta)$-PAC algorithm that runs in $\mathrm{poly}(n,k,1/\epsilon,\log(1/\delta))$ time, makes $\mathrm{poly}(n)$ calls to an ERM oracle for $\mathcal{F}$, and has a sample complexity of
@@ 1.3 Related Work
Most closely related to our work are the previous studies of Collaborative PAC Learning [BHPQ17, NZ18, CZZ18, Qia18] and related fields such as multi-task learning [HK22], multi-distribution learning [HJZ22, AHZ23, Pen23, ZZC+23], federated learning [MMR+17, MSS19, CCD23] and multi-source domain adaptation [MMR08, KL19, MMR+21].
In multi-task learning/multi-source domain adaptation, there are $n$ distributions, each with a fixed number of samples, and our goal is to use these samples to learn a hypothesis that has a small risk on some target distribution. A line of works [BDBC+10, KL19, MMR+21] studied the generalization risk in this scenario, but their bounds all depend on the discrepancy among $n$ distributions and contain non-vanishing residual constants. To avoid this residual constant in the bounds, [HK22] considered a Bernstein condition assumption on the hypothesis class and some transferrability assumptions between the $n$ source distributions and the target distribution. They studied the minimax rate of this learning scenario, and gave a nearly-optimal adaptation algorithm. Specially, some works studied the multi-task linear regression [YZW+20, HXL+23]. [YZW+20] considered linear regression from multiple distributions, where all tasks share the same input feature covariate $\mathbf{X}\in\mathbb{R}^{m\times d}$, but with different labeling functions. They designed the Hard Parameter Sharing estimator and established an excess risk upper bound of $O(\frac{d\sigma^{2}}{m}+\mathrm{heterogeneity})$. Recently, [HXL+23] proposed the $s$-sparse heterogeneity assumption among the labeling functions in multi-task linear regression, and designed an algorithm which achieves an $O(\frac{s\sigma^{2}}{m_{i}}+\frac{d\sigma^{2}}{\sum_{i=1}^{n}m_{i}})$ excess risk on the $i$-th task. Notice that when $s$ is smaller than $d$, this bound is strictly sharper than individual learning bound $O(\frac{d\sigma^{2}}{m_{i}})$. The difference between our scenario and multi-task learning is that we allow the learner to draw an arbitrary number of samples from each distribution, instead of assuming that each distribution only has a fixed number of samples.
@@
Multi-distribution learning was recently proposed by [HJZ22], where $n$ players try to learn a single model $\hat{f}$, that can have an $\epsilon$ excess error on the worst case distribution among $n$ players, i.e., $\max_{i\in[n]}\mathrm{err}_{\mathcal{D}_{i}}(\hat{f})\leq\epsilon+\mathrm{OPT}$. They gave an algorithm with sample complexity $O(\frac{d\log n}{\epsilon^{2}}+\frac{nd\log(d/\epsilon)}{\epsilon})$, and proved a lower bound of $\tilde{\Omega}(\frac{d+k}{\epsilon^{2}})$. Note that this is a more pessimistic learning guarantee than ours, since the value $\mathrm{OPT}$ can be very large. Very recently, two concurrent papers [Pen23, ZZC+23] gave algorithms that match this lower bound, resolving some of the open problems formulated in [AHZ23].
@@
Another recent line of work [KSS+20, KSKO20, DJKS23, JSK+23] studied learning mixtures of linear regressions from data batches. In these setups, there are $k$ unknown linear regression models. Each data batch consists of labeled examples produced by one of the linear models chosen randomly. This line of work gave trade-offs between the number of batches and the batch size in order for the parameters of the $k$ linear models to be efficiently learnable.
In comparison, our model allows the learner to adaptively sample from the data distributions, whereas the batch sizes are fixed in the model of learning with batches. We also note that the results for learning with batches require assumptions on the marginal distribution, such as Gaussianity or certain hypercontractivity and condition number properties. Also, except the recent work of [JSK+23], they all required the marginal distribution of the instance to be the same across all batches.
@@ 3 Sample Complexity Upper Bound
Lemma 11 implies that in general, $\mathcal{F}_{G,k}$ has a VC dimension of $O(kd+|G|\log k)$. To gain some intuition behind the bound in Lemma 11, suppose that $\mathcal{F}$ is a finite class of size $2^{d}$. By Definition 2, the size of $\mathcal{F}_{n,k}$ is at most $|\mathcal{F}|^{k}\cdot k^{n}=2^{kd+n\log_{2}k}$, and the bound in the lemma immediately follows. The actual proof, of course, needs to deal with the case that $\mathcal{F}$ is larger or even infinite.
The lemma improves a previous result of [CM12, Theorem 1], which upper bounds the VC dimension of $\mathcal{F}_{n,k}$ (called the class of “hard $k$-shared task classifiers”) by $O(kd\log(nkd)+n\log k)$. Note that this bound can be looser than the one in Lemma 11 by a logarithmic factor.
@@ 5.2 Hardness under Two Labeling Functions
We first reduce the general subset sum problem to a special case, in which the $n$ numbers sum up to $2^{n+1}$ and the target value is exactly $2^{n}$. Let ($(\{a_{i}\}_{i\in[m]},t)$ be an instance of subset sum (i.e., deciding whether there exists $T\subseteq[m]$ such that $\sum_{i\in T}a_{i}=t$). Let $s=\sum_{i=1}^{m}a_{i}$ and pick $n=\max\{m+2,\lfloor\log_{2}s\rfloor+1\}$ such that $n-m\geq 2$ and $2^{n}>s$. We pad $n-m$ numbers to the instance, such that $a_{m+1}=a_{m+2}=\cdots=a_{n-2}=0$, $a_{n-1}=2^{n}-t$, $a_{n}=2^{n}-(s-t)$. Note that the numbers now sum up to
@@
Also note that the size of the instance increases at most polynomially after the padding: A natural representation of the original subset sum instance ($(\{a_{i}\}_{i\in[m]},t)$ takes at least $m+\log_{2}s$ bits. In the new instance, there are $n=O(m+\log s)$ numbers that sum up to $2^{n+1}$, so its representation takes at most $O(n^{2})$ bits, which is at most quadratic in the size of the original instance.
@@ 6 An Efficient Algorithm for Identical Marginals
In this section, we prove Theorem 4, which addresses the special case that $\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{n}$ share the same marginal distribution over $\mathcal{X}$. In this case, we show that there is a simple algorithm that efficiently clusters the distributions and learn accurate classifiers using $\ll nd$ samples. The algorithm is formally defined as Algorithm 1, and follows a similar approach to the lifelong learning algorithms of [BBV15, PU16].
@@ 7.2 Technical Lemmas
The proof is based on similar techniques to the analysis of [Qia18] for a different variant of collaborative learning, in which a small fraction of the data sources are adversarial.